# Conquering the Communication Constraints to Enable Large Pre-Trained Models in Federated Learning

### v3

@@ IV-C RQ2: Capability Analysis
To study and understand our second research question (RQ2, Section IV), we analyze the impact of the domain gap between the model pre-training dataset and the dataset for FL (Section IV-C) and systemically perform experiments on CIFAR-100 across different federated learning scenarios by varying client status and data distribution (Section IV-C).
Capability with Domain Gap. Domain gap is a realistic concern when deploying pre-trained models for downstream tasks. As shown in TABLES I and II, the performance on different datasets varies a lot. Besides the difference in the difficulty between all datasets, the domain gap to the pre-trained dataset is also a key concern here. To further discuss the impact of the domain gap for each method, we use our default setting shown in Section IV-A on the image domain as an instance to visualize and quantify the domain gap between each downstream dataset and the pre-training dataset. Specifically, we adapt Linear Discriminant Analysis (LDA) for all extracted features for the test samples in each dataset from the pre-trained backbone to reduce the dimension. We compute the center of each dataset and then compute the distance to the center of the pre-training dataset as the quantifying result of the domain gap, as shown in Fig. 4a.
@@
TABLE IV: Robustness analysis for privacy-preserving. The red number indicates the performance difference under DP.
@@
Capability with Different FL Settings. In application scenarios, the setting of federated learning can vary substantially. It is important to show the capability to maintain high performance in diverse settings. We present results for all approaches with different client availability ratios and data distributions in TABLE III and draw the following conclusions from the experiments:
First, we see that fine-tuning the pre-trained model shows a significant improvement over training from scratch, especially in heterogeneous scenarios. This finding is in agreement with other very recent works [8, 7], which note the stabilization effect of pre-trained initialization in federated optimization. When only fine-tuning the head, the performance is still much better than training the entire model from scratch (Scratch) but remains low in comparison to other methods across all settings. We again find that head-tuning simply lacks adaptation ability, holding too closely to the upstream representation.
IV-D RQ3: Robustness Analysis
@@
TABLE V: Robustness analysis for data scarcity. $K$ indicates the total sample number of all clients.
@@
To integrate DP, we apply a Gaussian mechanism within the local optimization of each iteration [68] with $\varepsilon=5$ and $\delta=0.001$. We maintain the remaining FL settings as described in Section IV-A, and show the results in TABLE IV. Interestingly, when comparing all methods, full fine-tuning experiences the sharpest drop with DP. This causes its accuracy to fall lower than all the FedPEFT prototypes. To understand this effect, we note that DP applies noise to all trainable parameter gradients. Full fine-tuning, therefore, requires such noise on all model parameters, resulting in a more pronounced negative effect on final performance. On the other hand, the other fine-tuning methods maintain some part of the backbone frozen and have significantly fewer trainable parameters on which adding noise is necessary, limiting the performance drop. Overall, FedPEFT allows for stronger accuracy in DP-enabled federated systems than even full fine-tuning while still maintaining extremely low communication needs.
Data Scarcity. We explore another common yet challenging robustness condition in FL; that is when very little data is available on individual clients. Such data scarcity scenarios are even a tricky problem in centralized training. For example, [69] show that fewer training data will incur damage to the pre-trained representation due to overfitting. In our evaluation for FL, we reduce the total sample number $K$ to $1000$, $1500$, and $2000$. As shown in TABLE V, we find that FedPEFT outperforms full fine-tuning and head-tuning under such low-data scenarios, further revealing its capability to appropriately adapt pre-trained representations to the FL task at hand.
@@
TABLE VI: Ablation study on network backbone and pre-training dataset. 1K and 21K mean pre-trained on ImageNet1K and ImageNet21K respectively. The server accuracy (acc) on CIFAR-100 is reported.
Method Backbone Pre-training Server Acc (%)
Full Fine-tuning ViT-B Supervised (21K) 92.09
ViT-B Supervised (1K) 92.09
ViT-S Supervised (21K) 89.33
ViT-B DINO (1K) 85.62
Head-tuning ViT-B Supervised (21K) 72.55
ViT-B Supervised (1K) 80.78
ViT-S Supervised (21K) 74.40
ViT-B DINO (1K) 77.17
FedPEFT-Bias ViT-B ImageNet-21K 91.02
ViT-B Supervised (1K) 90.73
ViT-S Supervised (21K) 88.40
ViT-B DINO (1K) 85.34
FedPEFT-Adapter ViT-B ImageNet-21K 87.99
ViT-B Supervised (1K) 88.69
ViT-S Supervised (21K) 84.15
ViT-B DINO (1K) 83.29
FedPEFT-Prompt ViT-B ImageNet-21K 89.63
ViT-B Supervised (1K) 89.29
ViT-S Supervised (21K) 87.39
ViT-B DINO (1K) 84.25
TABLE VII: Ablation study on network backbone and pre-training dataset. The server accuracy on UCF-101 is reported.
Method Backbone Pre-training Server Acc (%)
Full Fine-tuning ViT-B Supervised (K400) 94.22
ViT-B Supervised (SSV2) 93.94
ViT-S Supervised (K400) 93.03
Head-tuning ViT-B Supervised (K400) 88.57
ViT-B Supervised (SSV2) 88.12
ViT-S Supervised (K400) 87.64
FedPEFT-Bias ViT-B Supervised (K400) 92.34
ViT-B Supervised (SSV2) 92.16
ViT-S Supervised (K400) 92.10
FedPEFT-Adapter ViT-B Supervised (K400) 92.82
ViT-B Supervised (SSV2) 91.97
ViT-S Supervised (K400) 91.48
FedPEFT-Prompt ViT-B Supervised (K400) 93.82
ViT-B Supervised (SSV2) 93.22
ViT-S Supervised (K400) 91.66
TABLE VIII: Ablation study on different federated learning algorithms. The server accuracy on CIFAR-100 and UCF-101 is reported for the image and video domain, respectively.
Method Image Video
FedAVG FedProx MOON FedAVG FedProx MOON
Full Fine-tuning 92.09 92.11 91.99 94.22 94.44 94.30
Head-tuning 72.55 72.55 72.55 88.57 88.57 88.57
FedPEFT-Bias 91.02 91.02 91.15 92.34 92.30 92.34
FedPEFT-Adapter 88.05 88.19 88.13 92.82 92.96 93.08
FedPEFT-Prompt 89.90 90.20 90.25 93.82 93.92 93.98
TABLE IX: Compatibility of FedPEFT with other PEFT methods. The server accuracy on CIFAR-100 and UCF-101 is reported for the image and video domain, respectively.
Method # Trainable Params Image Video
Full Fine-tuning 85.88M 92.09 94.22
Head-tuning 0.08M 72.55 88.57
FedPEFT-Bias 0.18M 91.02 92.34
FedPEFT-Adapter 0.23M 88.05 92.82
FedPEFT-Prompt 0.17M 89.90 93.82
FedPEFT-Prefix [34] 0.08M 86.26 87.68
FedPEFT-LoRA [29] 0.22M 91.06 92.54
@@ IV-E Insights and Takeaways
  • The effectiveness of FedPEFT is particularly pronounced in real-world applications characterized by stringent privacy requirements or data scarcity. Even under such conditions, FedPEFT-Bias and FedPEFT-Prompt maintain exceptional performance across image and video domains, surpassing other methods.
V Discussion
V-A Ablation Study
In the experiments above, we maintain a consistent backbone (ViT-B). However, since model size requirements of different federated systems may vary, we further investigate the use of ViT-S with FedPEFT and all baselines. Additionally, we study the effect of pre-training dataset size and paradigm on the final performance of all approaches. CIFAR-100 is used for the experiments.
Impact of Model Size. We replace ViT-B with a smaller version, ViT-S. As expected, most methods experience a slight performance drop. Nonetheless, FedPEFT methods maintain strong performance even with smaller models. Surprisingly, the performance of head-tuning is not damaged by reducing the model size but actually increases slightly. We interpret it as due to the fixed representation of head-tuning. Representations from a pre-trained model can have similar semantics whatever the model size, thus the performance of head-tuning is stable.
Impact of Pre-training Dataset. We also evaluate the performance when a backbone is trained from ImageNet-21K and ImageNet-1K [70]. Interestingly, the performance stays relatively consistent, with Head-tuning and FedPEFT-Adapter achieving higher accuracy with the ImageNet-1K models. This phenomenon indicates that the performance of fine-tuning is not always proportional to the data size, but overall FedPEFT performance is consistent with different scales of the pre-training dataset.
Impact of Pre-training Paradigm. Additionally, we evaluate one self-supervised pre-trained backbone: DINO [71] for the image domain. FedPEFT still achieves comparable results with full fine-tuning and outperforms Head-tuning, showing the robustness of FedPEFT.
Impact of Federated Learning Algorithm. Besides the widely-used FedAvg algorithm, FedPEFT focuses on the local training and communication stages, thus being orthogonal and compatible with other algorithms. We evaluate FedPEFT along with the baseline with FedProx [2] and MOON [21] in TABLE VIII and find the performance can be further improved when more advanced algorithms are integrated.
V-B Extension Capability
In this paper, we mainly focus on the computer vision fields with three base FedPEFT prototypes. We further show FedPEFT’s extension capability to be extended with other PEFT methods and for natural language processing tasks.
PEFT Methods. LoRA [29] and Prefix [34] are another popular backbone and input adaptation methods. As shown in TABLE IX, FedPEFT is compatible, and the performance of FedPEFT can be further improved with LoRA, while the communication costs can be further reduced with Prefix.
NLP task. We evaluate the performance of FedPEFT and baselines in text classification in the AG NEWS [72] dataset under the setting $M=64,N=8,\alpha=1$ with MiniBERT [73] as the pre-trained backbone in TABLE X. FedPEFT-Bias, with the highest fine-tuning freedom, still provides robust performance, which is consistent with our insights in CV. Further, a follow-up work of FedPEFT gives a study in NLP, further indicating the extension capability of FedPEFT [42].
TABLE X: Performance of FedPEFT on NLP task. # TP means the number of trainable parameters, and FedPEFT-B, -A, and -P mean FedPEFT prototypes with bias, adapter, and prompt.
Method Full Head FedPEFT-B FedPEFT-A FedPEFT-P
# TP 11M 1.0K 12.8K 17.7K 31.7K
Server Acc. 88.68 46.97 83.20 77.80 44.82
VI Conclusion
@@
VII Acknowledgement
This work is partially supported by the NSF/Intel Partnership on MLWiNS under Grant No. 2003198.

### v5

@@ IV-C RQ2: Capability Analysis
To study and understand our second research question (RQ2, Section IV), we analyze the impact of the domain gap between the model pre-training dataset and the dataset for FL (Section 4) and systemically perform experiments on CIFAR-100 across different federated learning scenarios by varying client status and data distribution (Section IV-C).
Figure 4: Visualization and analysis of domain gap.
Capability with Domain Gap. Domain gap is a realistic concern when deploying pre-trained models for downstream tasks. As shown in Tab. I, the performance on different datasets varies a lot. Besides the difference in the difficulty between all datasets, the domain gap to the pre-trained dataset is also a key concern here since Resisc45 is a remote sensing dataset, and PCam is a medical dataset. To further discuss the impact of the domain gap for each method, we use our default setting shown in Section IV-A on the image domain as an instance to visualize and quantify the domain gap between each downstream dataset and the pre-training dataset. Specifically, we adapt Linear Discriminant Analysis (LDA) for all extracted features for the test samples in each dataset from the pre-trained backbone to reduce the dimension. We compute the center of each dataset and then compute the distance to the center of the pre-training dataset as the quantifying result of the domain gap, as shown in Fig. 4a.
@@
Capability with Different FL Settings. In application scenarios, the setting of federated learning can vary substantially. It is important to show the capability to maintain high performance in diverse settings. We present results for all approaches with different client availability ratios and data distributions in Table II and draw the following conclusions from the experiments:
First, we see that fine-tuning the pre-trained model shows a significant improvement over training from scratch, especially in heterogeneous scenarios. This finding is in agreement with other very recent works [8, 7], which note the stabilization effect of pre-trained initialization in federated optimization. When only fine-tuning the head, the performance is still much better than training the entire model from scratch but remains low in comparison to other methods across all settings. We again find that head-tuning simply lacks adaptation ability, holding too closely to the upstream representation.
IV-D RQ3: Robustness Analysis
TABLE III: Robustness analysis for privacy-preserving. The red number indicates the performance difference.
@@ IV-D RQ3: Robustness Analysis
To integrate DP, we apply a Gaussian mechanism within the local optimization of each iteration [66] with $\varepsilon=5$ and $\delta=0.001$. We maintain the remaining FL settings as described in Section IV-A, and show the results in Table III. Interestingly, when comparing all methods, full fine-tuning experiences the sharpest drop with DP. This causes its accuracy to fall lower than all the FedPEFT prototypes. To understand this effect, we note that DP applies noise to all trainable parameter gradients. Full fine-tuning, therefore, requires such noise on all model parameters, resulting in a more pronounced negative effect on final performance. On the other hand, the other fine-tuning methods maintain some part of the backbone frozen and have significantly fewer trainable parameters on which adding noise is necessary, limiting the performance drop. Overall, FedPEFT allows for stronger accuracy in DP-enabled federated systems than even full fine-tuning while still maintaining extremely low communication needs.
TABLE IV: Robustness analysis for data scarcity. $K$ indicates the total sample number of all clients.
@@
Data Scarcity. We explore another common yet challenging robustness condition in FL; that is when very little data is available on individual clients. Such data scarcity scenarios are even a tricky problem in centralized training. Fewer training data will incur damage to the pre-trained representation due to overfitting. In our evaluation for FL, we reduce the total sample number $K$ to $1000$, $1500$, and $2000$. As shown in Table IV, we find that FedPEFT outperforms full fine-tuning and head-tuning under such low-data scenarios, further revealing its capability to appropriately adapt pre-trained representations to the FL task at hand.
@@ IV-E Insights and Takeaways
  • The effectiveness of FedPEFT is further illustrated in real-world applications characterized by stringent privacy requirements or data scarcity. Even under such conditions, FedPEFT-Bias and FedPEFT-Prompt maintain exceptional performance across image and video domains.
V Conclusion
@@
VI Acknowledgement
This work is partially supported by the NSF/Intel Partnership on MLWiNS under Grant No. 2003198 and the NSF Grant No. 2008447.